# DART-LLM: Dependency-Aware Multi-Robot Task Decomposition and Execution using Large Language Models

### v1

@@ IV-A Experimental Setup
To evaluate the performance of the DART-LLM system across different task sets, this study uses construction robots, which are highly sensitive to the order of sub-tasks, as test subjects to assess the performance of DART-LLM. Errors in the sequence of sub-tasks may lead to experimental failure; for instance, if a dump truck departs the soil pile before the excavator completes unloading, the loading and unloading operation will fail. This study designed a test set comprising three task levels (L1, L2, L3) and six test scenarios. Each task level includes different tasks to evaluate the performance of various models in a simulated environment. The task numbering reflects its respective level.
The experimental environment is based on the Unity simulation platform. The simulator uses the PhysX physics engine. The simulation environment includes two C30R crawler transport robots produced by Japan’s Yanmar and one ZX120 excavator produced by Japan’s Hitachi Construction Machinery.
Table IV lists the specific tasks, required skills, and applicable robot types for each task level. Level 1 tasks mainly assess the ability of a single robot to perform simple tasks; Level 2 tasks require multiple robots to work cooperatively; Level 3 tasks are more complex, requiring robot teams to consider dependencies and the sequence of subtasks.
@@
Level 1 L1-T1 Inspect a puddle N1, N2, N3, N4 Excavator or Dump Truck
L1-T2 Clear an obstacle N1, N2, N3, N4, FE1, FE2 Excavator
Level 2 L2-T1 Excavate soil N1, N2, N3, N4, FE1, FD1, FE2, FD2 Excavator, Dump Truck
L2-T2 Transport soil to the dump truck’s initial position N1, N2, N3, N4, FE1, FD1, FE2, FD2 Excavator, Dump Truck
Level 3 L3-T1 Clear the obstacle, then dig soil N1, N2, N3, N4, FE1, FD1, FE2, FD2 Excavator, Dump Truck
L3-T2 Clear the obstacle, then inspect the puddle N1, N2, N3, N4, FE1, FE2, FE2, FD2 Excavator, Dump Truck
@@ IV-B Evaluation Metrics
To comprehensively assess the performance of different models in the tasks mentioned above, this study uses the following key metrics:
@@ IV-C Experimental Results
IV-C1 Testing on the Test Set Using Different LLMs
This study tested DART-LLM using GPT-4o, GPT-3.5-turbo, Llama3.1, and Claude3.5. Table V shows the evaluation results of different models on tasks of L1, L2, and L3. The results indicate that for task level L1, all models achieved perfect scores across all metrics (SR=1.00, IPA=1.00, DSR=1.00, SGSR=1.00). As task complexity increases (from L1 to L3), performance differences among models become apparent. In task level L2, the IPA scores for all models were 1. The SR of the Llama 3.1 model was 0.87, higher than GPT-3.5-turbo’s 0.83, influenced by SGSR. This may be attributed to Llama 3.1’s better instruction-following ability, resulting in the generation of compliant JSON commands. For GPT-3.5-turbo, the SR dropped to 0.83 in L2 tasks, with a DSR of 0.87 and SGSR of 0.83. In contrast, the GPT-4o model maintained a high level of performance with an SR of 0.97, and both DSR and SGSR scores of 0.97. This can be attributed to GPT-4o’s enhanced reasoning and instruction-following abilities, allowing it to think more logically and generate compliant commands, leading to higher DSR and SGSR scores. The Claude 3.5 model performed worse than GPT-4o but outperformed both the Llama 3.1 and GPT-3.5-turbo models. All models performed well in terms of IPA, which can be attributed to the introduction of the structured JSON format and dependency-aware task decomposition method, allowing models to better focus on the task decomposition itself. At task level L3 (the highest complexity), the GPT-4o model continued to excel, maintaining scores similar to those in task level L2, with an SR of 0.97, IPA of 1.00, and both DSR and SGSR scores of 0.97. Other models showed a decline in performance in L3 tasks, particularly GPT-3.5-turbo, whose SR dropped to 0.75 due to its insufficient instruction-following capability. Overall, the DART-LLM system showed optimal performance when using the GPT-4o model, especially in handling more complex tasks. Notably, the Llama 3.1 model, with only 8B parameters, performed better than GPT-3.5-turbo, despite not achieving the best results. In task level L3, while Llama 3.1’s DSR score was lower than that of GPT-3.5-turbo, its higher SGSR resulted in an SR that was superior to GPT-3.5-turbo. This indicates that using smaller models with higher instruction-following capabilities is the best choice for the DART-LLM system, as smaller models offer better deployability and real-time performance.
TABLE V: Evaluation results of different models on L1, L2, and L3 tasks
@@ IV-C1 Testing on the Test Set Using Different LLMs
SR IPA DSR SGSR SR IPA DSR SGSR SR IPA DSR SGSR
@@
1.00 1.00 1.00 1.00 0.89 1.00 0.89 0.97 0.89 1.00 0.89 0.93
@@ IV-C2 Ablation Study
To investigate the impact of dependencies on the efficiency of multi-robot systems, we utilized the Llama3.1 model to test six task sets under conditions with and without the dependency-aware approach. The experimental results, as shown in Fig. 2, indicate that the dependency-aware approach significantly saves time in complex tasks. This is because the dependency-aware approach allows subtasks to be executed sequentially while also enabling parallel execution of subtasks without dependency relationships, indicating that the use of the dependency-aware approach enhances the system’s logical capabilities.
Figure 2: Comparison of Dependency-Aware and Non-Dependency-Aware Approaches across Various Tasks.
@@ IV-C3 Deployment in Real World
To evaluate the deployability of DART-LLM, we tested it in a real-world setting. Fig. 3 shows the results of the L2-T1 task. The results indicate that DART-LLM demonstrates good applicability on actual robots. For more experiments, please visit the project website project-dart-llm.
Figure 3: DART-LLM (Llama3.1) in the L2-T1 Task Using Real Robots: The sequence begins at 0 [s] with the initial positioning of the Dump Truck at the starting location. At 8 [s], the Excavator prepares to dig soil. By 24 [s], the Excavator transfers the excavated soil into the Dump Truck, completing the loading operation. At 48 [s], the loaded Dump Truck navigates towards the designated dumping area. At 56 [s], the Dump Truck unloads the soil at the puddle location. Finally, at 64 [s], the Dump Truck completes the task by retracting its vessel.
@@ V Conclusion and Future Work
This paper presents the DART-LLM, which leverages LLMs to achieve dependency-aware task decomposition and execution in multi-robot systems. By decomposing tasks into multiple subtasks with dependencies, DART-LLM effectively manages complex task sequences, facilitating parallel execution and collaborative cooperation in multi-robot systems. Experimental results validate the effectiveness of DART-LLM in handling complex tasks and dependencies, with the system showing optimal performance when using the GPT-4o model. Importantly, even when using smaller models like Llama 3.1 8B, DART-LLM demonstrated high success rates and efficiency, proving the system’s adaptability and robustness across different model sizes. Future work includes enhancing DART-LLM’s applicability to larger-scale robot teams.
@@ ACKNOWLEDGMENT
This work is supported by JST [Moonshot Research and Development], Grant Number [JPMJMS2032]. This work was submitted to the 2025 IEEE International Conference on Robotics and Automation on September 15, 2024.

### v2

@@ IV-A Experimental Setup
To evaluate the performance of the DART-LLM system across different task sets and conduct a quantitative comparison with baseline methods, this study employs construction robots that are highly sensitive to subtask sequencing. These robots serve as test subjects to evaluate the system’s effectiveness. Errors in subtask sequencing can lead to task failure. For instance, if a dump truck departs from the soil pile before the excavator has completed unloading, the loading and unloading operation will be unsuccessful. We constructed a benchmark dataset designed specifically for evaluating natural language-driven task planning in multi-robot construction scenarios. The dataset includes three task levels: L1, L2, and L3. The task numbering reflects their respective levels. L1 tasks are fundamental tasks involving a single robot equipped with all necessary skills, eliminating the need for multi-robot coordination. L2 tasks require collaboration among multiple robots. These tasks consist of subtasks that must be executed in a specific sequence, though each subtask can still be completed independently by a single robot. L3 tasks are complex, requiring multiple robots to work together. These tasks are decomposed into multiple interdependent subtasks, which must follow a strict execution order but may allow some parallel execution where feasible. The dataset comprises 102 high-level instructions: 47 L1 tasks, 33 L2 tasks, and 22 L3 tasks. It is available on our project website: project-dart-llm. Task execution can be visually validated in both simulated and physical robot environments. The simulation environment is built on the Unity platform and employs the PhysX physics engine. Both the simulation and physical environments feature two C30R tracked transport robots manufactured by Yanmar (Japan) and one ZX120 excavator manufactured by Hitachi Construction Machinery (Japan). Table IV presents two specific example tasks for each task level, along with the required skills and applicable robot types.
@@
Level 1 L1-T1-001 Inspect a puddle N1, N2, N3, N4 Excavator or Dump Truck
L1-T2-001 Clear an obstacle N1, N2, N3, N4, FE1, FE2 Excavator
Level 2 L2-T1-001 Excavate soil N1, N2, N3, N4, FE1, FD1, FE2, FD2 Excavator, Dump Truck
L2-T2-001 Transport soil to the dump truck’s initial position N1, N2, N3, N4, FE1, FD1, FE2, FD2 Excavator, Dump Truck
Level 3 L3-T1-001 Clear the obstacle, then dig soil N1, N2, N3, N4, FE1, FD1, FE2, FD2 Excavator, Dump Truck
L3-T2-001 Clear the obstacle, then inspect the puddle N1, N2, N3, N4, FE1, FE2, FE2, FD2 Excavator, Dump Truck
@@ IV-B Evaluation Metrics
To evaluate the performance of the proposed approach on the dataset, this study employs five evaluation metrics:
@@
  • RTR (Response Time Reliability): Assesses the model’s response time stability across multiple executions, using average response time and standard deviation to evaluate performance consistency under different conditions.
The evaluation relies on dataset ground truth to measure each metric.
@@ IV-C Experimental Results
IV-C1 Evaluation Against Baselines Using Different LLMs
This study evaluated DART-LLM using multiple foundation models: Llama-3.1-8B [17], GPT-4o [18], GPT-3.5-turbo [19], Claude-3.5-Haiku [20], and DeepSeek-r1-671B [21]. We also compared our approach with the SMART-LLM baseline implemented with Llama-3.1-8B and DeepSeek-r1-671B. Table V shows the evaluation results across tasks of L1, L2, and L3 complexity levels. Notably, all 102 test cases were unseen data, using few-shot examples solely for prompting and excluding them from the test set. All evaluation metrics range from 0 to 1, with higher values indicating better performance.
The results indicate that for task level L1, all implementations achieved perfect scores across SR, IPA, DSR, and SGSR metrics (all 1.00). However, notable differences appear in the RTR metric, where DART-LLM with Llama3.1 achieved the highest score of 0.96, significantly outperforming other models including SMART-LLM implementations (0.24 and 0.08).
As task complexity increases to L2, performance differences become more pronounced. DART-LLM with DeepSeek-r1 achieved the highest SR of 0.97. GPT-4o followed with 0.96, Claude3.5 with 0.90, GPT-3.5-turbo with 0.87, and Llama3.1 with 0.85. All DART-LLM implementations maintained perfect IPA scores, demonstrating the effectiveness of proposed structured JSON format and dependency-aware task decomposition method. In contrast, the SMART-LLM baseline showed substantial performance degradation, with SR scores of 0.36 and 0.78 for Llama3.1 and DeepSeek-r1 respectively, and reduced IPA scores (0.57 and 0.78). The explicit specification of dependencies in proposed approach significantly improved model logical reasoning capabilities, as evidenced by the high DSR scores across all DART-LLM implementations compared to the SMART-LLM baseline. For RTR, DART-LLM with Llama3.1 again demonstrated superior performance (0.90) compared to other models, while SMART-LLM implementations struggled significantly (0.19 and 0.04).
At task level L3, DART-LLM with DeepSeek-r1 demonstrated the best performance with an SR of 0.94. GPT-4o followed with 0.93, while Claude3.5, Llama3.1, and GPT-3.5-turbo achieved 0.86, 0.84, and 0.80 respectively. The SMART-LLM baseline performance degraded substantially, with SR scores dropping to 0.24 and 0.65 for Llama3.1 and DeepSeek-r1 respectively.
The structured JSON format consistently contributed to high RTR scores for DART-LLM (Llama3.1), maintaining 0.96, 0.90, and 0.86 across L1, L2, and L3 respectively. In contrast, the SMART-LLM baseline implementations showed markedly lower RTR scores (ranging from 0.02 to 0.24), primarily due to its direct generation of more complex Python code, which led to increased planning time for code generation. While the SMART-LLM baseline achieved reasonable SR scores with DeepSeek-r1, its low RTR indicates significant practical limitations in real-world applications where code execution is essential.
Overall, the DART-LLM system showed superior performance when using DeepSeek-r1 for SR metrics and Llama3.1 for RTR metrics. Notably, the Llama3.1 model, with only 8B parameters, performed competitively with much larger models and achieved significantly higher RTR scores than all other approaches. This suggests that using smaller models with proposed JSON-based command generation approach provides an excellent balance between performance and deployability, offering better real-time capabilities while maintaining high success rates.
TABLE V: Evaluation of DART-LLM and SMART-LLM baseline across L1, L2, and L3 tasks.
@@ IV-C1 Evaluation Against Baselines Using Different LLMs
SR IPA DSR SGSR RTR SR IPA DSR SGSR RTR SR IPA DSR SGSR RTR
@@
1.00 1.00 1.00 1.00 0.70 0.90 1.00 0.90 0.95 0.63 0.86 1.00 0.86 0.90 0.60
DART-LLM (DeepSeek-r1)
1.00 1.00 1.00 1.00 0.30 0.97 1.00 0.97 0.97 0.25 0.94 1.00 0.94 0.94 0.20
SMART-LLM(Llama3.1) [12]
1.00 1.00 1.00 1.00 0.24 0.36 0.57 0.46 0.36 0.19 0.24 0.37 0.43 0.24 0.16
SMART-LLM(DeepSeek-r1) [12]
1.00 1.00 1.00 1.00 0.08 0.78 0.78 0.91 0.87 0.04 0.65 0.65 0.86 0.78 0.02
@@ IV-C2 Ablation Study
To evaluate the effectiveness of proposed dependency-aware approach, we conducted an ablation study on L3 complexity tasks using five different LLMs. We compared two conditions: ”With Dependency” (i.e., using a DAG to represent task dependencies) and ”Without Dependency” (i.e., not using a DAG to specify task relationships).
Fig. 3 presents the SR under both conditions. The experimental results indicate that the ”Without Dependency” condition (i.e., without using a DAG) significantly reduces the performance of all models. Among them, Llama3.1 exhibited the largest performance drop (from 0.84 to 0.45). Even state-of-the-art models such as GPT-4o and DeepSeek-r1 experienced notable declines (from 0.93 to 0.85 and from 0.94 to 0.89, respectively). The pronounced performance gap in smaller models suggests that explicit dependency modeling via DAG can effectively compensate for their limited reasoning capabilities, demonstrating that proposed approach enables smaller, more deployable models to achieve competitive performance on complex tasks.
Figure 3: Success Rate (SR) comparison between ”With Dependencies” (using DAG) and ”Without Dependencies” (not using DAG) for L3 complexity tasks across different models.
@@ IV-C3 Deployment in Real World
To evaluate the deployment capability of DART-LLM, we conducted tests in a real-world environment. Fig. 4 presents the results of the L2-T1-001 task using the Llama3.1 model. The results indicate that DART-LLM demonstrates good applicability on real robots.
For more experiments, please visit the project website project-dart-llm.
Figure 4: DART-LLM (Llama3.1) in the L2-T1-001 Task Using Real Robots: The sequence begins at 0 [s] with the initial positioning of the Dump Truck at the starting location. At 8 [s], the Excavator prepares to dig soil. By 24 [s], the Excavator transfers the excavated soil into the Dump Truck, completing the loading operation. At 48 [s], the loaded Dump Truck navigates towards the designated dumping area. At 56 [s], the Dump Truck unloads the soil at the puddle location. Finally, at 64 [s], the Dump Truck completes the task by retracting its vessel.
@@ V Conclusion and Future Work
This paper presented DART-LLM, a novel framework that leverages LLMs and DAG-based dependency modeling for coordinated multi-robot task execution. The proposed system integrates four key modules: a QA LLM module for instruction parsing and dependency-aware task decomposition, a Breakdown Function module for task parsing and robot assignment, an Actuation module for executing robot-specific skills, and a VLM-based object detector module for environmental perception. This architecture demonstrated superior performance compared to existing methods across all evaluation metrics, with DART-LLM (DeepSeek-r1) achieving a 94% success rate on complex tasks while maintaining perfect instruction parsing accuracy. The smaller Llama3.1 model exhibited exceptional response time reliability, making it suitable for resource-constrained deployments. Furthermore, ablation experiments reveal that DAG-based dependency modeling significantly enhances model performance. Notably, it compensates for the limited reasoning capabilities of smaller models, enabling them to execute complex tasks more effectively. Real-world testing validated the system’s practical applicability. Future work will focus on scaling DART-LLM for larger multi-robot teams to enhance system applicability. Additionally, optimizing the trade-off between model size and performance will be investigated to accommodate diverse deployment scenarios.
@@ ACKNOWLEDGMENT
This work is supported by JST [Moonshot Research and Development], Grant Number [JPMJMS2032]. The work was first submitted to an IEEE conference on September 15, 2024.